OpenStudy (bananas1234):

what is the approximate distance between the points (5, -6, 6) and (-4, 5, -3)

OpenStudy (rayep12):

can u give us a graph or a picture?

OpenStudy (bananas1234):

I was not given one, sorry

OpenStudy (rayep12):

k

satellite73 (satellite73):

\[\sqrt{(x_1-x_2)^2+(y_1-y_2)^2+(z_1-z_2)^2}\] just ilke in two dimensions

OpenStudy (bananas1234):

can you explain it to me step by step. Im having trouble understanding it.

OpenStudy (bananas1234):

(5 + 4) Would be the first part right?

satellite73 (satellite73):

\[\sqrt{(5+4)^2+...}\]

satellite73 (satellite73):

\[\sqrt{(5+4)^2+(-6-5)^2+(6+3)^2}\] and a calculator

satellite73 (satellite73):

\[\sqrt{9^2+11^2+9^2}\] whatever that is

OpenStudy (bananas1234):

how would i solve that part?

OpenStudy (bananas1234):

@satellite73

satellite73 (satellite73):

with a calculator

OpenStudy (bananas1234):

but that gives me a huge number

satellite73 (satellite73):

not me

satellite73 (satellite73):

it give me about 16.8